# Comparing Bayesian and Frequentist Inference in Biological Models: A Comparative Analysis of Accuracy, Uncertainty, and Identifiability

### v1

@@ A.1 LV-Model (BFF) Histograms
Figure S1: Histograms for predator and prey data of $(\alpha,\beta,\delta,\gamma)$ Parameters
@@
Figure S2: Histograms for predator data only of $(\alpha,\beta,\delta,\gamma)$ Parameters
@@
Figure S3: Histograms for prey data only of $(\alpha,\beta,\delta,\gamma)$ Parameters
A.2 LV-model (QDF) Histograms
@@
Figure S4: Histograms for predator-prey data of $(\alpha,\beta,\delta,\gamma)$ Parameters
@@
Figure S5: Histograms for predator data only of $(\alpha,\beta,\delta,\gamma)$ Parameters
@@
A.3 GLM-Model-Lung Injury BayesianFitForecast (BFF) Histograms
Figure S7: Histograms for GLM data of r, k, and p Parameters
A.4 GLM-Model-Lung Injury (QDF) Histograms
Figure S8: Histograms r,p, and k Parameters
A.5 GLM-Model-Mpox (BFF) Histograms
Figure S9: Histograms for GLM data of r, k, and p Parameters
A.6 GLM-Model-Mpox (QDF) Histograms
Figure S10: Histograms for GLM data of r, k, and p Parameters
A.7 SEIUR-Model (BFF) Histograms
Figure S11: Histograms ($\beta_{0}$, $\beta_{1}$, $q_{1}$, $\rho$, $\kappa$, and $\gamma_{1}$ ) Parameters
A.8 SEIUR-Model (QDF) Histograms
Figure S12: Histograms for ($\beta_{0}$, $\beta_{1}$, $q_{1}$, $\rho$, $\kappa$, and $\gamma_{1}$ ) Parameters

### v2

@@ 7.1 LV-Model (BFF) Histograms
Figure 20: Histograms of the Lotka–Volterra parameters $(\alpha,\beta,\delta,\gamma)$.
@@
Figure 21: Histograms for predator data only of $(\alpha,\beta,\delta,\gamma)$ Parameters
@@
Figure 22: Histograms for prey data only of $(\alpha,\beta,\delta,\gamma)$ Parameters
7.2 LV-model (QDF) Histograms
@@
Figure 23: Histograms for predator-prey data of $(\alpha,\beta,\delta,\gamma)$ Parameters
@@
Figure 24: Histograms for predator data only of $(\alpha,\beta,\delta,\gamma)$ Parameters
@@
7.3 GLM-Model-Lung Injury (BFF) Histograms
Figure 26: Histograms for GLM data of r, k, and p Parameters
7.4 GLM-Model-Lung Injury (QDF) Histograms
Figure 27: Histograms r,p, and k Parameters
7.5 GLM-Model-Mpox (BFF) Histograms
Figure 28: Histograms for GLM data of r, k, and p Parameters
7.6 GLM-Model-Mpox (QDF) Histograms
Figure 29: Histograms for GLM data of r, k, and p Parameters
7.7 SEIUR-Model (BFF) Histograms
Figure 30: Histograms ($\beta_{0}$, $\beta_{1}$, $q$, rhorho, kappakappa, and $sigma_{1}$ ) Parameters
7.8 SEIUR-Model (QDF) Histograms
Figure 31: Histograms for ($\beta_{0}$, $\beta_{1}$, $q$, $\rho$, $\kappa$, and $\gamma_{1}$ ) Parameters